# MedMimic: A Physician-Inspired Multimodal Fusion Framework for Early Diagnosing Fever of Unknown Origin

### v1

@@ Abstract
Fever of unknown origin (FUO) poses a significant diagnostic challenge, often requiring extensive evaluations. While recent efforts integrating clinical features with 18F-FDG PET/CT imaging have improved diagnostic accuracy, conventional feature extraction methods based on manually defined features often lack robustness and generalizability. Furthermore, the scarcity of annotated 18F-FDG PET/CT datasets limits the effective training of deep learning models.To overcome these challenges, we propose Medical Mimicry (MedMimic), a framework inspired by real-world diagnostic processes. MedMimic leverages pre-trained models—including DINOv2, Vision Transformer, and ResNet-18—to transform high-dimensional 18F-FDG PET/CT imaging data into low-dimensional, semantically meaningful feature tensors. A learnable self-attention-based fusion network then integrates these imaging features with clinical data, producing compact yet discriminative representations for downstream classification. Our dataset was collected from Sichuan University West China Hospital, one of China’s top-tier medical institutions, and comprised 607 consecutive FUO patients who underwent 18F-FDG PET/CT. After excluding cases that did not meet the inclusion criteria, 416 patient admissions from January 2017 to December 2023 were included in the study. Our proposed multimodal fusion classification network (MFCN) achieved macro-averaged area under the receiver operating characteristic curve (macro-AUROC) values ranging from 0.8654 to 0.9291 across all seven decomposed tasks, outperforming competing machine learning (ML) and single-modality deep learning (DL) methods. Additionally, ablation studies validated the effectiveness of the proposed MFCN model. Furthermore, we employed a five-fold cross-validation strategy to ensure the robustness of our results. By integrating the strengths of pre-trained large models and deep learning, MedMimic offers a novel perspective and an effective solution for disease classification.
@@ 4.3.1 Image Feature Extraction
As illustrated in Fig. 4, we extract image features from PET/CT scans using four distinct methods: PCA[35], ResNet-18, ViT, and DINOv2. Each method generates a slice-level feature vector of dimension $b_{k}$ for every image slice. In PCA, each slice is flattened, and principal component analysis is applied to reduce dimensionality. With ResNet-18, the slice is processed through the network, and the output from the global average pooling layer is used as the feature representation. In ViT, each slice is partitioned into patches and processed by a Vision Transformer to generate an embedding. DINOv2 follows a similar Transformer-based architecture but incorporates a specialized self-supervised pretraining strategy for enhanced feature learning.
@@
As shown in Fig. 5, Principal Component Analysis (PCA) serves as a baseline for comparison with advanced pre-trained models such as ResNet-18, ViT, and DINOv2. Each CT slice, denoted by $\mathbf{X}^{CT}$, is treated as a single-channel $n\times n$ square image. First, each image is flattened into a one-dimensional vector by concatenating its pixel values sequentially. These vectors are assembled into a data matrix, where each row corresponds to a CT slice, and each column represents a specific pixel position. To correct for global intensity variations, the mean value computed at each pixel position across all slices is subtracted, resulting in a zero-centered dataset. Next, the covariance matrix is computed to capture pixel intensity relationships. Singular value decomposition (SVD) is applied, and the top $b_{1}$ eigenvectors associated with the largest eigenvalues are retained, forming a projection matrix that transforms the original high-dimensional data into a lower-dimensional space. This reduces the number of features while preserving dominant variance for subsequent modeling tasks.
@@ 4.3.2 Multimodal Fusion Classification Network
5:  Initialization:
6:   Initialize convolution network parameters ConvK, ConvQ, ConvV;
7:  Step1: Tensor Alignment
8:  Normalize feature tensors:
9:   $\text{Norm}(\mathcal{F}^{CT})$, $\text{Norm}(\mathcal{F}^{PET})$;
10:  Apply zero-padding for consistency:
11:   $\mathcal{F}^{masked}\leftarrow\text{Pad}(\mathcal{F}^{CT},\mathcal{F}^{PET},\mathcal{Z})$;
12:  Fuse normalized clinical and masked features:
13:   $\mathcal{F}^{fused}\leftarrow\text{Concat}(\text{Expand}(\text{Norm}(\mathbf{A})),\mathcal{F}^{masked})$;
14:  Step2: Learnable Self-Attention Based Recalibration
15:  Apply $1\times 1$ convolution and reshape the tensor:
16:   $\mathbf{K}\leftarrow$ Reshape(ConvK($\mathcal{F}^{fused}$));
17:   $\mathbf{Q}\leftarrow$ Reshape(ConvQ($\mathcal{F}^{fused}$));
18:   $\mathbf{V}\leftarrow$ Reshape(ConvV($\mathcal{F}^{fused}$));
19:  Compute the attention map:
20:   $\mathbf{\alpha}\leftarrow\text{SoftMax}\left(\frac{\mathbf{K}^{T}\mathbf{Q}}{\sqrt{I}}\right)$;
21:  Compute the attention tensor and reshape:
22:   $\mathbf{F}^{attn}\leftarrow\text{Reshape}(\mathbf{V}\mathbf{\alpha})$;
23:  Perform global average pooling along $M_{\max}$ dimension:
24:   $\mathbf{F}^{fused}\leftarrow\text{AvgPool}(\mathcal{F}^{attn})$;
25:  Step3: Optimization
26:   Compute model loss via Eq. (2);
27:   Update ConvK, ConvQ, ConvV via gradient descent;
28:  Output: Fused feature matrix $\mathbf{F}^{fuse}$.
@@ 5.4 Ablation Studies
To evaluate the contribution of different components within the MFCN model, we designed and tested seven experimental configurations: (1) without residual connections, (2) without attention mechanisms, (3) without dropout, and (4) variations in the ResNet architecture with 3, 4, 5, or 6 layers. To ensure optimal feature representation, we utilized the DINOv2 model for semantic feature extraction from CT/PET images and conducted experiments using three data modalities: Clinical Data, CT, and PET.
@@
Furthermore, the results highlight the critical role of dropout, learnable attention layers, and residual connections in achieving high-quality outcomes. Dropout mitigates overfitting by randomly deactivating neurons during training, introducing redundancy, and enhancing generalization. Learnable attention layers assign adaptive weights to information from different modalities, improving the model’s ability to capture long-range dependencies and salient features. Residual connections, through skip connections, preserve original input information, mitigate vanishing gradient issues in small-sample datasets, and enhance the stability and efficiency of deep network training. Additionally, increasing the number of ResNet layers further improves performance by enabling the model to learn more complex feature hierarchies. A deeper architecture facilitates the extraction of richer representations, leading to superior performance on challenging tasks and greater robustness when generalizing to new data.

### v2

@@ Abstract
Fever of unknown origin (FUO) presents a major diagnostic challenge, often necessitating extensive evaluations. Although integrating clinical features with 18F-FDG PET/CT imaging has enhanced diagnostic accuracy, conventional feature extraction can lack generalizability, and limited annotated datasets restrict deep learning. To address these issues, we propose Medical Mimicry (MedMimic), a framework that leverages pre-trained models—including DINOv2, Vision Transformer, and ResNet-18—to transform high-dimensional PET/CT data into semantically meaningful feature tensors. A learnable self-attention-based fusion network then integrates imaging features with clinical data, producing compact yet discriminative representations for downstream classification. Our dataset from Sichuan University West China Hospital included 607 consecutive FUO patients (January 2017 to December 2023), with 416 admissions remaining after exclusions. Specifically, we decomposed the diagnostic challenge into seven classification tasks reflecting common clinical etiologies, enabling more targeted model evaluation. Our multimodal fusion classification network (MFCN) achieved macro-averaged area under the receiver operating characteristic curve (macro-AUROC) values of 0.8654 to 0.9291 across these tasks, outperforming competing machine learning (ML) and single-modality deep learning (DL) approaches. Ablation studies confirmed the efficacy of each MFCN component, while five-fold cross-validation demonstrated consistent performance. By harnessing the strengths of pre-trained large models and deep learning, MedMimic offers a novel perspective and practical solution for disease classification. These findings underscore the potential of leveraging large-scale, pre-trained architectures in complex clinical scenarios, bridging the gap between imaging data and clinical decision-making. Future work will explore external validations in multi-center settings and the extension of MedMimic to other diagnostic domains, emphasizing its versatility and clinical value.
@@ 4.3.1 Image Feature Extraction
As illustrated in Fig. 4, image features from PET/CT scans are extracted using four distinct methods: PCA [35], ResNet-18, ViT, and DINOv2. Each method generates a slice-level feature vector of dimension $b_{k}$ through $f_{pretrained}$ for every image slice. In PCA, each slice is flattened, and principal component analysis is applied to reduce dimensionality. With ResNet-18, the slice is processed through the network, and the output from the global average pooling layer serves as the feature representation. In ViT, each slice is partitioned into patches and processed by a Vision Transformer to generate an embedding. DINOv2 employs a similar Transformer-based architecture but integrates a specialized self-supervised pretraining strategy to enhance feature learning.
@@
As shown in Fig. 5, Principal Component Analysis (PCA) serves as a baseline for comparison with advanced pre-trained models such as ResNet-18, ViT, and DINOv2. Each CT slice, denoted by $\mathbf{X}^{CT}$, is treated as a single-channel $n\times n$ square image. First, each image is flattened into a one-dimensional vector by sequentially concatenating its pixel values. These vectors are assembled into a data matrix, where each row represents a CT slice and each column corresponds to a specific pixel position. To correct for global intensity variations, the mean intensity at each pixel position across all slices is subtracted, yielding a zero-centered dataset. The covariance matrix is then computed to capture pixel intensity relationships. Singular value decomposition (SVD) is applied, and the top $b_{1}$ eigenvectors corresponding to the largest eigenvalues are retained. These eigenvectors form a projection matrix that transforms the original high-dimensional data into a lower-dimensional space, reducing the number of features while preserving dominant variance for subsequent modeling tasks.
@@ 4.3.2 Multimodal Fusion Classification Network
5:   Ground-truth label for each patient: $\mathbf{y}$;
6:  Initialization:
7:   Initialize convolution network parameters ConvK, ConvQ, ConvV;
8:  Step 1: Tensor Alignment
9:  Normalize feature tensors:
10:   $\text{Norm}(\mathcal{F}^{CT})$, $\text{Norm}(\mathcal{F}^{PET})$;
11:  Apply zero-padding for consistency:
12:   $\mathcal{F}^{masked}\leftarrow\text{Pad}(\mathcal{F}^{CT},\mathcal{F}^{PET},\mathcal{Z})$;
13:  Fuse normalized clinical and masked features:
14:   $\mathcal{F}^{fused}\leftarrow\text{Concat}(\text{Expand}(\text{Norm}(\mathbf{A})),\mathcal{F}^{masked})$;
15:  Step 2: Learnable Self-Attention Based Recalibration
16:  Apply $1\times 1$ convolution and reshape the tensor:
17:   $\mathbf{K}\leftarrow$ Reshape(ConvK($\mathcal{F}^{fused}$));
18:   $\mathbf{Q}\leftarrow$ Reshape(ConvQ($\mathcal{F}^{fused}$));
19:   $\mathbf{V}\leftarrow$ Reshape(ConvV($\mathcal{F}^{fused}$));
20:  Compute the attention map:
21:   $\mathbf{\alpha}\leftarrow\text{SoftMax}\left(\frac{\mathbf{K}^{T}\mathbf{Q}}{\sqrt{I}}\right)$;
22:  Compute the attention tensor and reshape:
23:   $\mathbf{F}^{attn}\leftarrow\text{Reshape}(\mathbf{V}\mathbf{\alpha})$;
24:  Perform global average pooling along $M_{\max}$ dimension:
25:   $\mathbf{F}^{fused}\leftarrow\text{AvgPool}(\mathcal{F}^{attn})$;
26:  Step 3: Optimization
27:   Compute model loss via Eq. (2);
28:   Update ConvK, ConvQ, ConvV via gradient descent;
29:  Output: Fused feature matrix $\mathbf{F}^{fuse}$.
@@ 5.4 Ablation Studies
To evaluate the contribution of different components within the MFCN model, we designed and tested seven experimental configurations: (1) without residual connections, (2) without attention mechanisms, (3) without dropout, (4) without convolutional layers, and (5) variations in the ResNet architecture with 3, 4, 5, or 6 layers. To ensure optimal feature representation, we utilized the DINOv2 model for semantic feature extraction from CT/PET images and conducted experiments using three data modalities: Clinical Data, CT, and PET.
@@
Furthermore, the results underscore the critical role of dropout, learnable attention layers, and residual connections in achieving high-quality outcomes. Dropout prevents overfitting by randomly deactivating neurons during training, introducing redundancy, and enhancing generalization. Learnable attention layers assign adaptive weights to information from different modalities, improving the model’s capacity to capture long-range dependencies and salient features, while convolutional layers facilitate the learnability of attention. Residual connections preserve original input information through skip connections, mitigate vanishing gradient issues in small-sample datasets, and enhance the stability and efficiency of deep network training. Additionally, increasing the number of ResNet layers further improves performance by enabling the model to learn more complex feature hierarchies. A deeper architecture facilitates the extraction of richer representations, leading to superior performance on challenging tasks and greater robustness in generalizing to new data.